OpenStudy (anonymous):

lim (x-4 ) / (sqareroot x)-2 x->4

OpenStudy (anonymous):

\[\lim_{x \rightarrow 4} \frac{ x-4 }{ \sqrt{x}-2 }\] In proper notation. How do I resolve this, I know it should equal 4, but trying to resolve the radical made this harder than easier.

OpenStudy (ash2326):

\[\lim_{x \rightarrow 4} \frac{ x-4 }{ \sqrt{x}-2 }\times \frac{\sqrt x+2}{\sqrt x +2}\] Can you try @Astrobuoy

OpenStudy (anonymous):

Yeah I was trying that earlier I get \[\frac{ x^{2/3}+2x-4\sqrt{x}-8 }{ x-4 }\]

OpenStudy (anonymous):

I F.O.I.L'd the numerator after doing the denominator.

OpenStudy (ash2326):

Just multiply the denominator terms, you'd get an awesome term

OpenStudy (anonymous):

So do you mean: Keep the top the same, x-4 times squareroot x+2, cancel the x-4 with the denom and them left with squareroot x+2?

OpenStudy (anonymous):

ahh! thus squareroot 4 which is 2 + 2 = 4, the solution! correct>?

OpenStudy (ash2326):

Absolutely :D

OpenStudy (anonymous):

I kept thinking the numerator had to be finished before proceeding. :\

OpenStudy (ash2326):

It happens sometimes :)

OpenStudy (anonymous):

Cheers! :)